OpenStudy (anonymous):

Identify whether the given value of the variable makes the equation true. 2k – 4.5 = 6.5; k = 5.5 A. true B. false

OpenStudy (anonymous):

all u hv to do is put value of k in equation so 2(5.5)-4.5 = 11-4.5 = 6.5= RHS so it is true

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

1 medal

OpenStudy (anonymous):

I am asking another Question, could you help me with that?